a man can row at 5 km/h in still water.he rows upstream 3 km back to his starting point.if the total time is 1 hour 15 min ,the speed of the current is
Answer by mananth(16946)

boat speed = 5 kmph
current speed = x kmph

upstream speed = 5 - x
downstream speed = 5 + x

Total distance = 3 km Total Time= 1.25hours
3 /( 5 - x )+ 3 /( 5 + x )= 1.25
3 ( 5 + x )+ 3 ( 5 - x )= 1.25
15 + 3 x + 15 -3 x = 1.25 ( 15^2-x^2)
30 = 1.25 ( 25 - x^2 )
30 = 31.25 -1.25 X^2
X ^2 = 1.25
X^2 1
x= 1 km/h
m.ananth@hotmail.ca